### v1

@@ 5.1.1 Performance of Various Models on Full Datasets
# 1 22.68 7.83 10.48 13.79 16.72 11.40 11.28 10.36 16.05 13.39 15.22 15.70 21.11 15.06 14.76 12.07 3
# 2 16.11 6.29 8.32 15.35 16.86 11.44 10.81 8.56 14.89 12.04 12.21 12.22 14.94 10.75 12.40 12.00 6
# 3 19.72 5.64 9.20 24.22 23.60 16.58 21.93 10.47 20.92 11.24 15.76 16.07 23.10 17.60 20.48 10.10 2
# 5 16.45 6.50 7.74 22.66 21.04 16.93 19.20 10.92 17.88 10.75 18.29 16.58 20.90 15.65 19.49 11.70 2
# 6 17.67 7.97 9.27 17.20 17.70 13.74 26.78 12.10 23.78 14.47 15.03 14.88 19.08 15.64 13.31 11.40 0
# 7 14.02 6.45 8.85 17.14 19.98 12.49 12.19 10.83 12.64 13.95 13.88 12.67 13.03 13.51 13.89 11.67 3
5.2 Mean Relative Gain
5.2.1 Mean Relative Gain of Models
To assess the comprehensive performance of each model, we calculate the mean relative gain for each model across all instructions on each dataset, as detailed in Eq. 3. The results are tabulated in Table 5. The metrics “Wins1” and “Wins3” denote the number of tasks where a model or an instruction achieves the highest and top-three aggregated performances, respectively. Similar to the best model performance metric, models equipped with the Flan-T5-XXL backbone, such as InstructBLIP, BLIP-2, and Flan-T5-XXL itself, exhibit strong performance. Closed-source models also show excellent results.
5.2.2 Mean Relative Gain of Instructions
We continue to evaluate the overall efficacy of each instruction by calculating the mean relative gain (MRGI) across all models, as detailed in Eq. 4. The outcomes are presented in Table 6. Due to the different evaluation datasets for LLMs and MLLMs—with LLMs excluding PuzzleVQA and MMMU and MLLMs including these datasets—we compute MRGI separately for each group. Our analysis reveals that Instructions # 1, # 2, # 5, and # 7 generally perform better, with their detailed designs illustrated in Figures 3 and 4. Notably, Instructions # 5 and # 7, which evolve from simple “Question-Answer” format of # 2, show superior performance. The effectiveness likely stems from the prior training of models on related QA tasks, where specific terms like “Question” and “Answer” enhance model responsiveness. Furthermore, the inclusion of option terms in Instructions # 5 and # 7 appears to positively influence performance on specific datasets.
5.3 Stability
5.3.1 Stability of Models
We exclusively use Eq. 5 to calculate stability metrics for models demonstrating strong overall performance, such as ChatGPT, Flan-T5-XXL, Gemini-V, BLIP2, and InstructBLIP, with the results displayed in Table 7. In terms of stability among open-source models, the Flan-T5 series exhibits minimal fluctuations across different instructions, indicating a lower sensitivity to instruction variability. Compared to BLIP2, InstructBLIP enhances stability by incorporating instruction tuning, effectively reducing the sensitivity of models to varying instructions in multimodal reasoning tasks with vision-text contexts. The improvement suggests that instruction tuning is a valuable strategy for enhancing model stability.
@@ 5.4 Adaptability
Models are trained with distinct pre-training settings, leading to varying preferences towards different instructions. We quantify the adaptability using the Global Top-K Hit Ratio, as detailed in Eq. 7. To thoroughly assess the adaptability of various models to different instructions across all full datasets, we have categorized all models into nine groups based on their language backbones. These groups include Closed-source, LLaMA LLMs, Other LLMs, MLLMs based on LLaMA, MLLMs based on Vicuna, MLLMs based on Flan-T5, LLaVA Series, LaVIT Series, and MLLMs based on other language backbones. The comprehensive results, covering all datasets, are illustrated in Figure 8. It’s important to note that evaluations for LLMs exclude the AlgoPuzzleVQA and MMMU datasets, while MLLMs include the two datasets.
@@ 5.4.2 Adaptability for Different Models
We advert that different models tend to perform better with specific instructions. (1) Closed-source models: Closed-source models exhibit their poorest performance on Instructions # 6 and # 9, suggesting that specific trigger words such as “AI:”, “Human:”, and “### Input” significantly impair their performance. (2) Same models with different version: The LLaMA model series, which includes LLaMA1-7B, LLaMA1-13B, LLaMA2-7B, and LLaMA2-13B, consistently shows strong performance on Instructions # 1 and # 2, indicating similar preferences across different models. Similarly, the LaVIN model branch excels with Instructions # 2 and # 7. In contrast, the LLaVA model branches demonstrate more varied preferences, as models of different sizes optimize performance using different instructions. The variation in the LLaVA models may be attributed to differences in training methodologies across different scales. (3) Various backbones of LLMs: We further categorize additional models into four groups based on large language model backbones. Related groups are listed that models using LLaMA, Vicuna, Flan-T5, and others as their LLMs. We note that models based on different backbones exhibit varied performance across diverse instructions. For instance, most models utilizing the LLaMA backbone excel on Instructions # 1, # 2, and # 7. However, certain models, such as OpenFlamingo and mPLUG-OWL, struggle with specific instructions, achieving an adaptability score of zero on # 4, among others. Models with Vicuna backbones generally underperform on Instructions # 8 and # 1. Conversely, the Flan-T5 series, which includes Flan-T5, BLIP-2, and InstructBLIP, shows consistent adaptability trends, performing well on Instructions # 9 and # 10 but less so on Instructions # 1 and # 8. Models leveraging other LLM backbones display optimal performance on a range of instructions, with Fromage, in particular, experiencing significant performance variability across different instructions.
@@ 6 Conclusion
We conduct comprehensive evaluations and establish benchmarks for 31 models, including 23 MLLMs, using 10 instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. Our framework, MM-InstructEval, introduces a range of multi-perspective metrics to assess these tasks comprehensively. These metrics include the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Through the comprehensive evaluation, we have drawn several important conclusions: (1) Models utilizing the Flan-T5-XXL backbone, which is based on an Encoder-Decoder architecture, consistently outperform those from the LLaMA series, which employ a Decoder-only architecture, in our multimodal reasoning evaluations. (2) Instructions formatted in a “Question-Answer” style tend to yield better performance. (3) There is a notable variation in model preferences for specific instructions across different models. (4) Excellent closed-source models, such as GPT-4V and Claude3, demonstrate superior overall performance across all tasks, excelling in challenging datasets. The study about Multimodal Large Language Models is a rapidly evolving field. Our research lays the groundwork for new exploratory directions and assesses the zero-shot performance of MLLMs in complex multimodal reasoning tasks. Looking forward, we aim to explore the performance of MLLMs in in-context learning scenarios in future studies.

### v2

@@ 5.3.1 Stability of Models
# 1 26.44 21.76 21.12 11.42 21.84 9.98 10.13 11.65 15.90 18.39 22.75 15.58 17.65 16.83 16.20 17.10 2
# 2 19.90 19.94 19.80 13.02 19.86 10.04 9.67 9.93 15.61 17.70 21.43 12.96 13.03 15.60 14.53 17.85 5
# 3 22.72 19.23 18.59 20.02 26.46 14.18 19.27 10.81 18.39 19.55 22.62 15.94 19.16 16.98 18.37 17.32 0
# 5 19.95 19.43 17.81 18.88 26.35 14.82 16.92 11.17 16.29 19.46 22.47 15.52 17.99 16.32 18.35 18.36 1
# 6 22.23 21.47 20.88 14.18 21.09 11.78 24.19 13.28 20.48 19.16 22.59 15.05 15.97 17.10 14.30 16.31 2
# 7 17.28 18.84 17.95 14.29 22.20 10.89 10.95 12.54 14.04 20.05 21.33 13.65 11.12 15.14 15.56 17.70 5
# 9 19.96 18.67 18.85 17.23 24.42 12.68 19.95 11.91 21.02 18.86 23.90 16.29 17.81 17.96 15.88 16.60 1
We exclusively use Eq. 5 to calculate stability metrics for models demonstrating strong overall performance, such as ChatGPT, Flan-T5-XXL, Gemini-V, LLaMA-3.2-11B-Vision, Qwen2-VL, Qwen2.5-VL, MiniCPM-V2.6, BLIP2, InstructBLIP, and others. The results are presented in Table 7. Among open-source models, the Flan-T5 series exhibits minimal fluctuations across different instructions, indicating lower sensitivity to instruction variability. Compared to BLIP2, InstructBLIP enhances stability by incorporating instruction tuning, effectively reducing the model’s sensitivity to varying instructions in multimodal reasoning tasks with vision-text contexts. This improvement highlights instruction tuning as a valuable strategy for enhancing model stability. Additionally, the recently released MiniCPM-V2.6 and Qwen2-VL-72B-Instruct combine exceptional performance with remarkably low sensitivity. This suggests that these models are specifically designed with instruction tuning and additional training in mind to optimize stability and performance.
@@ 5.4 Adaptability
Models are trained with distinct pre-training settings, leading to varying preferences towards different instructions. We quantify the adaptability using the Global Top-K Hit Ratio, as detailed in Eq. 7. To thoroughly assess the adaptability of various models to different instructions across all full datasets, we have categorized all models into nine groups based on their language backbones. These groups include Closed-source, LLaMA LLMs, Other LLMs, MLLMs based on LLaMA, MLLMs based on Vicuna, MLLMs based on Flan-T5, LLaVA Series, LaVIT Series, DeepSeek Series, Qwen-VL Series, and MLLMs based on other language backbones. The comprehensive results, covering all datasets, are illustrated in Figure 8. It’s important to note that evaluations for LLMs exclude the AlgoPuzzleVQA and MMMU datasets, while MLLMs include the two datasets.
@@ 5.4.2 Adaptability for Different Models
We advert that different models tend to perform better with specific instructions. (1) Closed-source models: Closed-source models exhibit their poorest performance on Instructions # 6 and # 9, suggesting that specific trigger words such as “AI:”, “Human:”, and “### Input” significantly impair their performance. (2) Same models with different version: The LLaMA model series, including LLaMA1-7B, LLaMA1-13B, LLaMA2-7B, LLaMA2-13B, and LLaMA-3.2-11B-Vision, consistently demonstrates strong performance on Instruction # 2, indicating similar preferences across different models. Similarly, the LaVIN model branch excels with Instructions # 2 and # 7. The Qwen-VL series achieves better performance on Instructions # 3, # 5, and # 7. In contrast, the LLaVA model branch and the DeepSeek series exhibit more varied preferences, as models of different sizes optimize performance using different instructions. This variation may be attributed to differences in training methodologies applied at different scales. (3) Various backbones of LLMs: We further categorize additional models into four groups based on large language model backbones. Related groups are listed that models using LLaMA, Vicuna, Flan-T5, and others as their LLMs. We note that models based on different backbones exhibit varied performance across diverse instructions. For instance, most models utilizing the LLaMA backbone excel on Instructions # 1, # 2, and # 7. However, certain models, such as OpenFlamingo and mPLUG-OWL, struggle with specific instructions, achieving an adaptability score of zero on # 4, among others. Models with Vicuna backbones generally underperform on Instructions # 8 and # 1. Conversely, the Flan-T5 series, which includes Flan-T5, BLIP-2, and InstructBLIP, shows consistent adaptability trends, performing well on Instructions # 9 and # 10 but less so on Instructions # 1 and # 8. Models leveraging other LLM backbones display optimal performance on a range of instructions, with Fromage, in particular, experiencing significant performance variability across different instructions.
@@ 6 Conclusion
In this work, we present MM-InstructEval, a comprehensive evaluation framework that assesses 45 models (including 36 MLLMs) across 16 diverse multimodal reasoning datasets using 10 different instruction formats. Our systematic evaluation framework introduces multiple innovative metrics, the Best Performance metric, Mean Relative Gain metric, Stability metric, and Adaptability metric, enabling a thorough and multi-faceted assessment of model capabilities. Through our extensive evaluation, we uncover several significant findings that provide valuable insights. First, newer models achieve comparable or superior performance with substantially fewer parameters compared to their predecessors. This finding indicates significant advances in model architecture design and training methodologies, suggesting a shift toward more efficient and practical implementations of MLLMs. Second, we reveal that “Question-Answer” style instructions generally lead to better performance across models. However, we also observe notable variations in instruction preferences across different models, highlighting the importance of careful prompt engineering in practical applications. Third, our evaluation demonstrates the superior capabilities of advanced closed-source models like GPT-4V and GLM-4V-plus, particularly in handling challenging datasets and complex reasoning tasks. This performance gap provides valuable benchmarks for ongoing development in open-source alternatives.
Research on multimodal large language models continues to advance at a rapid pace. Our comprehensive evaluation framework and findings provide a robust foundation for understanding these systems’ capabilities and limitations. This work bridges the gap between the theoretical understanding of multimodal reasoning and practical implementation considerations. Our future research agenda focuses on two key directions. First, we aim to enhance evaluation methodologies by developing more rigorous test cases and introducing holistic metrics that assess the generalization capabilities of models, ethical alignment, and performance consistency across diverse scenarios. Second, we will advance model performance through innovative approaches, particularly exploring contextual and reinforcement learning techniques to strengthen sophisticated reasoning capabilities across modalities.
Acknowledgements
Thanks to all co-authors for their hard work. The work is supported by the National Natural Science Foundation of China (62172086, 62272092), and the Fundamental Research Funds for the Central Universities under Grant (N25XQD004).